OpenStudy (anonymous):

Where are the x-intercepts for f(x) = 4 cos(2x − π) from x = 0 to x = 2π? x = 3 pi over 4, x = 5 pi over 4 x = pi over 2, x = 3 pi over 2 x = 3 pi over 4, x = 5 pi over 4, x = 7 pi over 4 x = pi over 2, x = 3 pi over 2, x = 5 pi over 2

OpenStudy (anonymous):

nobody has answered this yet i cant find the formula @phi do you know it?

OpenStudy (phi):

the x-intercepts means what is the x value when f(x) value is zero. in other words, find all the x values so that 4 cos(2x − π) = 0

OpenStudy (anonymous):

and then when they =2 aswell?

OpenStudy (phi):

you could simplify the function. for example divide both sides by 4 to get cos(2x − π) = 0/4 which is just cos(2x − π) = 0

OpenStudy (anonymous):

and that gets cos simplified

OpenStudy (phi):

***and then when they =2 as well?*** I don't understand this question. Please ask a different way

OpenStudy (anonymous):

ok.. u said i need to find when x=0.. i see that is an intercept in the question would i find it for when x is 2 because of the 2 intercept in the problem

OpenStudy (phi):

to answer the question cos(2x − π) = 0 you can either 1) solve for x (i.e. find the x's that work) or 2) test the choices given as possible answers. Use a calculator and be sure to be in radian mode.

OpenStudy (anonymous):

what exactly do you mean by radian mode?

OpenStudy (phi):

to solve, we can use cos(A-B)= cos(A)cos(B) + sin(A)sin(B) and write cos(2x-pi) = cos(2x)cos(pi) + sin(2x)sin(pi) you should(?) know sin(pi) is 0, and cosI(pi) is -1 so you get cos(2x-pi) = -cos(2x) setting equal to 0 -cos(2x)= 0 or multiply both sides by -1 to get just cos(2x)= 0 what angles is cos(y) zero? pi/2, pi/2 + pi, pi/2+2pi, etc set 2x equal to each of these, and solve for x